OpenStudy (brooke..help00):

WIL FAN AND MEDAL SO PLEASE HELP?? If g(x) = x2 + 2, find g(3). 9 8 11 6

OpenStudy (jiteshmeghwal9):

replace x by 3 in g(x)

OpenStudy (brooke..help00):

\[G \left( 3 \right)= X2\]

OpenStudy (jingle535):

Brookey replace all the x's with 3

OpenStudy (jiteshmeghwal9):

g(3)=3^2+2

OpenStudy (brooke..help00):

Oh- I'm smart. ok

OpenStudy (jingle535):

These can be confusing its ok :)

OpenStudy (brooke..help00):

\[g \left( 3 \right)= 11\]

OpenStudy (brooke..help00):

Right?

OpenStudy (jiteshmeghwal9):

yup